What volume of 0.67 mol/L sulphuric acid can be neutralized by 22.7 mL of 0.385 mol/L barium hydroxide
1
Expert's answer
2020-05-15T09:37:03-0400

H2SO4+Ba(OH)2=BaSO4+H2O


1.Find the amount of barium hydroxide substance:

"M=\frac{n}{V}"


"0.385=\frac{n}{0.0227}"


n=0.0087395


2.Find the volume of sulfuric acid:

"M=\frac{n}{V}"

"0.67=\frac{0.0087395}{V}"


V=0.013 L